A last minute grocery shopper uses a force of 125 N to push a 72 kg shopping cart down the 18.7 meter long baking aisle. The cart starts at one end of the aisle at a speed of 0.92 m/s and accelerates at a rate of 1.38 m/s2.

How fast is the cart going when it reaches the end of the aisle?
1
Expert's answer
2020-11-30T14:50:05-0500

"s=\frac{v^2-v_0^2}{2a}\to v=\sqrt{2as+v_0^2}=\sqrt{2\cdot 1.38\cdot 18.7+0.92^2}=7.2(m/s)"